Three resistances of 4Ξ©, 8Ξ© π‘Žπ‘›π‘‘ 16Ξ© are connected in a parallel to a 220 V supply. Calculate the equivalent resistance and current in circuit?


1
Expert's answer
2020-10-20T07:10:24-0400


The equivalent resistance "R" of a parallel combination of resistors "R_1, R_2,R_3" is given by

"\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}"

Given "R_1=4\Omega, R_2=8\Omega, R_3=16\Omega"

"\therefore" The equivalent resistance of the given circuit is

"\frac{1}{R}=\frac{1}{4\Omega}+\frac{1}{8\Omega}+\frac{1}{16\Omega}\\
\Rightarrow \frac{1}{R}=\frac{4+2+1}{16\Omega}\\
\Rightarrow \frac{1}{R}=\frac{7}{16\Omega}\\
\Rightarrow R=\frac{16}{7}\Omega\\
\Rightarrow R=2.286\Omega"

The current in the circuit is "I=\frac{V}{R}=\frac{220V}{2.286\Omega}=96.24A"


Answer: Equivalent resistance = "2.286\Omega" , current in the circuit = "96.24A"